OpenStudy (anonymous):

Simplify square root of negative 121

OpenStudy (jdoe0001):

well, what's \(\bf \sqrt{121} \quad ?\)

OpenStudy (anonymous):

its negative 121 and its 11 but itsays simplify

OpenStudy (anonymous):

lol its 11

OpenStudy (anonymous):

Take a pic i dont get it

OpenStudy (jdoe0001):

\(\bf -121 \implies 121 \times -1 \implies 11^2 \times -1\\\quad \\\quad \\ \sqrt{-121} \implies \sqrt{121 \times -1} \implies \sqrt{11^2 \times -1} \implies \sqrt{11^2} \times \sqrt{-1}\)